OpenStudy (anonymous):

Find the width of a rectangle if the length is 12 cm and the perimeter is 40 cm Find the width of a rectangle if the length is 12 cm and the perimeter is 40 cm @Mathematics

hero (hero):

(40-24)/2

jhonyy9 (jhonyy9):

us 2w+2l=P

hero (hero):

l = 8

OpenStudy (anonymous):

yea i got it thanx hero

OpenStudy (anonymous):

perimeter= Sum of all sides so, 40=2*(12+b); 20=12+b; b=8; njoy